The Treasury bill rate is 5% and the market risk premium is 8%.

Project Beta Internal Rate of Return, %
       P 1.00       16            
       Q 0       8            
       R 2.00     22            
       S 0.40       9            
       T 1.90       20            
a.

What are the project costs of capital for new ventures with betas of .75 and 1.55? (Do not round intermediate calculations. Round your answers to 2 decimal places.)

Beta       Cost of Capital
0.75           %   
  1.55           %   
b.

Which of the following capital investments have positive NPVs? (You may select more than one answer. Single click the box with the question mark to produce a check mark for a correct answer and double click the box with the question mark to empty the box for a wrong answer.)

  • Pchecked
  • Qunanswered
  • Tunanswered
  • Runanswered
  • Sunanswered
Answer #1
Cost of Capital = Risk free rate + beta*market risk premium
Project Beta Internal Rate of Return, % Cost of capital
       P 1 16             13 %
       Q 0 8             5 %
       R 2 22             21 %
       S 0.4 9             8.2 %
       T 1.9 20             20.2 %
Cost of capital for new ventures
Beta 0.75 = 5+0.75*8 = 11%
Beta 1.55 = 5+ 1.55*8 = 17.4%
NPV will positive for those projects whose IRR is higher than cost of capital
i.e. Projects P,Q,R and S